If 32.00 g of MgCO3 decomposes in this way, the mass of solid product expected is _______.
1
Expert's answer
2020-01-30T06:52:39-0500

MgCO3 = MgO +CO2

CO2 - gas

MgO - solid

mMgO = nMgO * MMgO

MMgO= 40 g/mol

nMgO = nMgCO3 = 32/ 84 = 0.38 mol

mMgO = 0.38* 40 = 15.24 g